# The Rise of Parameter Specialization for Knowledge Storage in Large Language Models

### v1

@@ 5.2 Finetuning Results
In addition to evaluating the model’s performance on SpecWiki’s Multi-choice Question and Open-ended Generation tests, as well as the Parameter Specialization scores, we also report two other metrics, Semantic Entropy (Kuhn et al., 2022) and Local Intrinsic Dimension (LID) (Yin et al., 2024), for measuring the extent of hallucination in the model’s output. These metrics help evaluate whether training strategies that enhance Parameter Specialization—by aligning better with the model’s knowledge retrieval mechanisms through a data-encoded strategy—can effectively reduce the unintended side effect of hallucination. For a detailed introduction to these two hallucination measurements, please refer to §C.3.
@@ 6 Conclusion
Appendix A Limitations and Future Work
In our work, we have only examined and validated knowledge parameter specialization within the MLP, and this was done by treating vectors in the MLP as units of analysis. However, at least this remains one of the knowledge storage methods that has been extensively validated so far (Geva et al., 2021a; Meng et al., 2022; Geva et al., 2023). In fact, knowledge may also reside within the attention module of transformer models (Geva et al., 2023).
@@ Appendix A Limitations and Future Work
In future work, we will progressively narrow the focus of our research to individual neurons in language models, aiming to measure and validate more precise Parameter Specialization and Parameter Superposition.
Appendix B Details of Dataset
B.1 Dataset Construction Prompts
@@ B.2 Manual Verification
B.3 Dataset Categories and Examples
@@
B.4 Validation of Popularity-Frequency Correlation
We included a simple experiment to validate the strong correlation between the popularity of concepts and their frequency in the pretraining data. To be specific, we used The Pile(Gao et al., 2020), which currently serves as a significant portion of the pretraining dataset for most large language models(Touvron et al., 2023a; Groeneveld et al., 2024; Team et al., 2024a), as an example of a pretraining corpus. We then counted the frequency with which each of the 525 concepts from SpecWiki appeared in all text segments of The Pile dataset via the Elasticsearch API(Elazar et al., 2024). Subsequently, we compared these frequencies with the popularity metrics for each concept and computed the corresponding Spearman’s rank correlation coefficient. The result is 0.814, indicating a strong correlation.
@@ B.4 Validation of Popularity-Frequency Correlation
Appendix C Details of Experiments
C.1 The implementation of the models
@@
C.2 Parameter Specialization Scores on OEG Setting
@@
C.3 Hallucination Metric Descriptions
In this experiment, we additionally incorporate two metrics, Semantic Entropy (Kuhn et al., 2022) and Local Intrinsic Dimension (LID) (Yin et al., 2024), to assess hallucination. This helps evaluate whether the finetuning methods that enhance Parameter Specialization also effectively mitigate the unintended side effect of hallucination in the model’s output.
@@ C.3 Hallucination Metric Descriptions
Semantic $\text{Semantic Entropy}=\frac{1}{|C|}\sum_{i=1}^{|C|}\log p(C_{i}|x)$
@@
The Local Intrinsic Dimension (LID) method detects hallucinations in Large Language Models by measuring the discrepancy in the local intrinsic dimension of model activations. This approach is grounded in the principle that LID represents the minimal number of activations required to characterize a data point, with truthful outputs exhibiting lower LID values due to their closer alignment with natural language structure, while hallucinated outputs tend to show higher LID values due to mixing human prompt and model distributions. Technically, the method employs Maximum Likelihood Estimation (MLE) using a Poisson process to approximate the count of neighbors surrounding sample points, computed through the formula $m(X_{i})=(1/(T-1)*\sum(log(Q_{T}/Q_{j})))^{-1}$, where T represents the number of nearest neighbors and $Q_{j}$ denotes the Euclidean distance to the j-th nearest neighbor. For more details about the Local Intrinsic Dimension metric, please refer to the work (Yin et al., 2024).

### v2

@@ 5.2 Finetuning Results
In addition to evaluating the model’s performance on SpecWiki’s Multi-choice Question and Open-ended Generation tests, as well as the Parameter Specialization scores, we also report two other metrics, Semantic Entropy (kuhn2022semantic) and Local Intrinsic Dimension (LID) (DBLP:conf/icml/YinSC24), for measuring the extent of hallucination in the model’s output. These metrics help evaluate whether training strategies that enhance Parameter Specialization—by aligning better with the model’s knowledge retrieval mechanisms through a data-encoded strategy—can effectively reduce the unintended side effect of hallucination. For a detailed introduction to these two hallucination measurements, please refer to §B.3.
@@ 6 Conclusion
7 Limitations and Future Work
In our work, we have only examined and validated knowledge parameter specialization within the MLP, and this was done by treating vectors in the MLP as units of analysis. However, at least this remains one of the knowledge storage methods that has been extensively validated so far (geva-etal-2021-transformer; meng2022locating; geva2023dissecting). In fact, knowledge may also reside within the attention module of transformer models (geva2023dissecting).
@@ 7 Limitations and Future Work
In future work, we will progressively narrow the focus of our research to individual neurons in language models, aiming to measure and validate more precise Parameter Specialization and Parameter Superposition. In addition, we will extend this concept to the study of other related model architectures, including Mixture of Experts (fedus2022switchtransformersscalingtrillion), which similarly enhances model performance by specializing expert parameters, as well as Sparse Auto-Encoders (huben2024sparse), which help clarify the model’s representations by leveraging a larger parameter space and mitigating the superposition of these features.
Acknowledgment
This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund (AcRF) Tier 1 grant, and funded through the SUTD Assistant Professorship Scheme (SAP 2025_001).
Appendix A Details of Dataset
A.1 Dataset Construction Prompts
@@ A.2 Manual Verification
A.3 Dataset Categories and Examples
@@
A.4 Validation of Popularity-Frequency Correlation
We included a simple experiment to validate the strong correlation between the popularity of concepts and their frequency in the pretraining data. To be specific, we used The Pile(gao2020pile800gbdatasetdiverse), which currently serves as a significant portion of the pretraining dataset for most large language models(touvron2023llamaopenefficientfoundation; groeneveld2024olmo; gemmateam2024gemmaopenmodelsbased), as an example of a pretraining corpus. We then counted the frequency with which each of the 525 concepts from SpecWiki appeared in all text segments of The Pile dataset via the Elasticsearch API(elazar2024whats). Subsequently, we compared these frequencies with the popularity metrics for each concept and computed the corresponding Spearman’s rank correlation coefficient. The result is 0.814, indicating a strong correlation.
@@ A.4 Validation of Popularity-Frequency Correlation
Appendix B Details of Experiments
B.1 The implementation of the models
@@
B.2 Parameter Specialization Scores on OEG Setting
@@
B.3 Hallucination Metric Descriptions
In this experiment, we additionally incorporate two metrics, Semantic Entropy (kuhn2022semantic) and Local Intrinsic Dimension (LID) (DBLP:conf/icml/YinSC24), to assess hallucination. This helps evaluate whether the finetuning methods that enhance Parameter Specialization also effectively mitigate the unintended side effect of hallucination in the model’s output.
@@ B.3 Hallucination Metric Descriptions
$\text{Semantic Entropy}=\frac{1}{|C|}\sum_{i=1}^{|C|}\log p(C_{i}|x)$
@@
The Local Intrinsic Dimension (LID) method detects hallucinations in Large Language Models by measuring the discrepancy in the local intrinsic dimension of model activations. This approach is grounded in the principle that LID represents the minimal number of activations required to characterize a data point, with truthful outputs exhibiting lower LID values due to their closer alignment with natural language structure, while hallucinated outputs tend to show higher LID values due to mixing human prompt and model distributions. Technically, the method employs Maximum Likelihood Estimation (MLE) using a Poisson process to approximate the count of neighbors surrounding sample points, computed through the formula $m(X_{i})=(1/(T-1)*\sum(log(Q_{T}/Q_{j})))^{-1}$, where T represents the number of nearest neighbors and $Q_{j}$ denotes the Euclidean distance to the j-th nearest neighbor. For more details about the Local Intrinsic Dimension metric, please refer to the work (DBLP:conf/icml/YinSC24).